the length of a rectangle is 8 inches more than its width.the perimeter is 80 inches. find the dimensions of the rectangle
Answer by ikleyn(52906)

.
L     + W = 40

(W+8) + W = 40


2W = 40 - 8 = 32  ====>  W = 32/2 = 16.


Answer.  The dimensions are 16 and 24 inches.